OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality x - 1 less than or greater to 7 or 2x greater than or equal to 22 A ) A number line with an open circle on 8, an open circle on 11, and shading in between. B ) A number line with an open circle on 8, shading to the left, and an open circle on 11, shading to the right. C) A number line with a closed circle on 8, shading to the left, and a closed circle on 11, shading to the right. D) A number line with a closed circle on 8, a closed circle on 11, and shading in between.

OpenStudy (anonymous):

What does less than or greater to mean? Is it supposed to be less than or equal to?

OpenStudy (anonymous):

less than

OpenStudy (anonymous):

What do you think it is?

OpenStudy (anonymous):

idk. thats why im asking

OpenStudy (anonymous):

But do you have a guess? If its greater than or equal to that means the circle will be shaded

OpenStudy (anonymous):

C and D are ruled out because the first part is just < so it wont be shaded, and would go to the left. The other is shaded and goes to the right.

OpenStudy (anonymous):

So its B

OpenStudy (anonymous):

x - 1 less than or greater to 7 or 2x greater than or equal to 22 \(\large x-1\le 7 \) or \(\large 2x \ge 22 \) is this the problem you need to solve?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so the wording on the first inequality is "less than or equal" not "less than or greater" (which does not make sense, btw...:) )

OpenStudy (anonymous):

solving these two inequalities should give: \(\large x\le 8 \) or \(\large x \ge 11 \) since these two inequalities include 8 and 11 the graphs should be closed circle. one is pointing to the left starting at 8 and the other pointing right starting at 11.

OpenStudy (anonymous):